# Multimodal Fact-Checking: An Agent-based Approach

### v1

@@ I Introduction
The proliferation of social media has intensified the spread of persuasive textual and visual misinformation, posing serious risks to public health and behavior [borges2022infodemics, wsj2025falseTariffHeadline, ARCURI2023106130]. Advances in generative AI further exacerbate this threat by enabling realistic, multimodal fake content [berman2024aimisinformation, reuters2025duterteDisinformation]. Traditional binary fact-checking is inadequate, as it lacks rationales and evidence that foster public trust [blanchar2024trump, Porter2021]. Narrow tasks like out-of-context (OOC) detection [qi2024sniffer] fall short in addressing real-world complexities. To enhance transparency and reliability, explainable multimodal fact-checking has become a crucial research focus.
@@ III-C Quality Control
The full prompt templates for both tasks are provided in the Appendix -A.
@@ IV-B Agents for Fact-Checking
As illustrated in Figure 4, the framework consists of five specialized agents: 1) Strategy Planning Agent (Agent-SP) , 2) Text Evidence Retrieval and Validation Agent (Agent-TR), 3) Image Retrieval and Analysis Agent (Agent-IR), 4) Reasoning Agent (Agent-R), 5) Explanation Generation Agent (Agent-EG). Each agent is designed with task-specific knowledge to mimic distinct components of real-world fact-checking workflows, and is deployed using prompt-engineering techniques with tool interaction. The simplified prompts are provided in the Appendix -B.
@@ V-B Implementation Details
Data Leakage Prevention: To ensure fair evaluation and prevent data leakage, we apply strict filtering to all retrieved evidence used before final reasoning for all external evidence based methods. Specifically, we exclude: (1) any evidence retrieved from Snopes or other major fact-checking websites (see details in Appendix -C), and (2) any evidence published after the original fact-checking date of the claim.
@@ V-D Performance on Explainability
so that higher values consistently indicate better performance. Full annotation guidelines and the interface are provided in the Appendix -E.
@@ V-E Ablation Study of Agents
We provide additional analysis on model performance across varying levels of sample difficulty in the Appendix -D.
@@ VI Conclusion
-A Dataset
Table S1 and Table S2 present the prompts used to filter samples in the RW-Post dataset, aiming to reduce label leakage from images and to exclude samples where only the text contributes to veracity assessment.
-B Prompts of Agents in AgentFact
-C Experimental Details: Fact-Checking Domain Filtering
To mitigate potential label leakage during evidence retrieval—particularly when external search engines return fact-checking articles that explicitly state the veracity of the claim—we applied a domain filtering strategy.
We excluded any evidence items whose source URLs contained substrings associated with known fact-checking organizations or services. This ensures that the model cannot trivially infer labels by relying on already-verified ground truth statements from professional sources. The filtering was implemented by checking whether the domain name or URL of the retrieved evidence contains any of the following substrings, which are associated with known fact-checking organizations or services. For brevity, a representative list is shown below; the complete list is available in our code repository.
snopes
politifact
factcheck
truthorfiction
hoax-slayer
eadstories
opensecrets
fullfact
checkyourfact
realitycheck
fact-check
This filter was applied consistently to all models that utilized external search (e.g., LEMMA, DEFAME, AgentFact) during evidence retrieval. By removing these high-authority verification sources, we ensure that the model’s performance more accurately reflects its ability to reason over raw evidence, rather than memorize or match against curated labels.
-D Experimental Results: Impact of Different Difficulty Levels.
To analyze model robustness under varying verification difficulty, we group claims based on the number of evidences referenced in their corresponding human-written fact-check articles. Importantly, these evidences are not provided to the model during inference. Instead, the evidence count is used solely as a proxy for the inherent difficulty of verifying a claim, reflecting the amount of effort required by human fact-checkers.
Figure S1 illustrates model performance across different difficulty levels for both binary (true vs. not_true) and three-class (true / false / unproven) verification settings. For binary evaluation, we collapse the original three-class labels into a two-class setting by treating true as the positive class and grouping all remaining labels (i.e., false and unproven) into a single negative class. This formulation reflects a practical fact-checking scenario where the primary objective is to determine whether a claim can be verified as true, while claims that are false or lack sufficient supporting evidence are both treated as not true.
Across both settings, we observe a clear non-monotonic performance trend with respect to claim difficulty. In particular, claims of moderate difficulty—those requiring 4 to 6 evidences in human fact-check articles—consistently yield the lowest Macro-F1 scores. This phenomenon is observed in both binary and three-class evaluations, indicating that such claims pose the greatest challenge to the model.
Interestingly, claims associated with very high evidence redundancy (9+ evidences) achieve the strongest performance in both evaluation settings. Despite being more complex in terms of information volume, these claims often correspond to highly visible events with strong cross-source consensus, which may facilitate model verification through information redundancy.
Overall, these results indicate that claim difficulty, as approximated by groundtruth evidence count, has a substantial impact on model robustness. Rather than exhibiting monotonic degradation, model performance follows a valley-shaped pattern, with moderate-difficulty claims being the most challenging. This highlights the importance of developing fact-checking models that are robust to fragmented and ambiguous evidence. We do not claim statistical significance for these differences, as the number of high-difficulty cases is relatively small.
-E Human evaluation
Figure S2 shows the interface used by human evaluators to score the outputs of different models, including the ground-truth results. The results of the three models (including the ground-truth) are randomly ordered for each sample, and the annotators are blind to which model produced each result.
The following guideline was given to human evaluators, outlining the definitions of all evaluation metrics and the corresponding rating criteria.
Task Definition
The goal of the evaluator (the annotator) is to assess the outputs produced by fact verification models. Each model aims to determine the authenticity of a given claim and provide corresponding explanations.
Each claim originates from a social media post. The post text and image are provided as contextual information to assist understanding, but the final evaluation focuses strictly on the claim itself.
Evaluation Interface Overview
In the annotation interface, each claim is presented together with its textual content, image context, and outputs generated by three different models. Each model output is shown as an independent card containing:
Authenticity judgment
Reasoning and explanation
Cited evidence
After finishing the evaluation for one claim, the results should be exported as a JSON file named according to the claim ID (e.g., 1.json, 2.json).
Each model output follows a structured JSON format:
"predicted_authenticity_class": "True",
"predicted_authenticity_reasons": "...",
"predicted_authenticity_key_points": [...],
"evidence_list": [...],
"image_analysis_result": [...],
"confidence_level": ""
Reasoning Hallucination
What to Evaluate
This dimension evaluates the factual soundness of the model’s reasoning. Specifically, check whether the reasoning provided in predicted_authenticity_reasons is directly supported by the evidence listed in evidence_list.
How to Judge
1.
Read the reasoning and key points.
2.
Identify factual claims such as numbers, dates, or causal statements.
3.
Verify whether these claims are supported by the cited evidence.
4.
If no concrete facts are invoked, evaluate logical coherence only.
Target Fields
predicted_authenticity_reasons
predicted_authenticity_key_points
| Score | Criteria |
| none | Reasoning is coherent and well-supported by evidence. |
| mild | Minor unsupported assumptions without affecting the main conclusion. |
| severe | Core reasoning is speculative or disconnected from evidence. |
Evidence Usage Hallucination
What to Evaluate
This dimension assesses whether the cited evidence is correctly used and accurately represented.
How to Judge
1.
Examine each item in evidence_list and image_analysis_result.
2.
Check whether the evidence supports the claims made.
3.
Identify any fabrication, misinterpretation, or irrelevance.
Target Fields
evidence_list
evidence_content_and_link
image_analysis_result
| Score | Criteria |
| none | Evidence is accurate, relevant, and supports the reasoning. |
| partial | Evidence is partially misused or weakly relevant. |
| full | Evidence is fabricated or contradicts the conclusion. |
Label Justification
What to Evaluate
This dimension examines whether the predicted authenticity label is justified.
How to Judge
1.
Identify the predicted label.
2.
Compare it with the reasoning and evidence.
3.
Check for overconfidence or contradiction.
Target Fields
predicted_authenticity_class
predicted_authenticity_reasons
| Score | Criteria |
| justified | Label is consistent with reasoning and evidence. |
| overconfident | Label is stronger than supported by evidence. |
| hallucinated | Label contradicts the reasoning or evidence. |

### v2

@@ I Introduction
The proliferation of social media has intensified the spread of persuasive textual and visual misinformation, posing serious risks to public health and behavior [borges2022infodemics, wsj2025falseTariffHeadline, ARCURI2023106130]. Advances in generative AI further exacerbate this threat by enabling realistic, multimodal fake content [berman2024aimisinformation, DeepfakeSurvey2024Wang, reuters2025duterteDisinformation]. Traditional binary fact-checking is inadequate, as it lacks rationales and evidence that foster public trust [blanchar2024trump, Porter2021]. Narrow tasks like out-of-context (OOC) detection [qi2024sniffer] and deepfake detection [NoiseDF2023Wang, wang2023dire] fall short in addressing real-world complexities. To enhance transparency and reliability, explainable multimodal fact-checking has become a crucial research focus.
@@ III-C Quality Control
The full prompt templates for both tasks are provided in the Appendix LABEL:app:dataset.
@@ IV-B Agents for Fact-Checking
As illustrated in Figure 4, the framework consists of five specialized agents: 1) Strategy Planning Agent (Agent-SP) , 2) Text Evidence Retrieval and Validation Agent (Agent-TR), 3) Image Retrieval and Analysis Agent (Agent-IR), 4) Reasoning Agent (Agent-R), 5) Explanation Generation Agent (Agent-EG). Each agent is designed with task-specific knowledge to mimic distinct components of real-world fact-checking workflows, and is deployed using prompt-engineering techniques with tool interaction. The simplified prompts are provided in the Appendix VI.
@@ V-B Implementation Details
Data Leakage Prevention: To ensure fair evaluation and prevent data leakage, we apply strict filtering to all retrieved evidence used before final reasoning for all external evidence based methods. Specifically, we exclude: (1) any evidence retrieved from Snopes or other major fact-checking websites (see details in Appendix VI), and (2) any evidence published after the original fact-checking date of the claim.
@@ V-D Performance on Explainability
so that higher values consistently indicate better performance. Full annotation guidelines and the interface are provided in the Appendix VI.
@@ V-E Ablation Study of Agents
We provide additional analysis on model performance across varying levels of sample difficulty in the Appendix VI.